OpenStudy (anonymous):

How to find the common multiple of 15 and 30?

myininaya (myininaya):

well 15*2=30 and 30*1=30

OpenStudy (anonymous):

3,5,15.

myininaya (myininaya):

common multiple is 30

OpenStudy (anonymous):

15 and 20 sorry.

myininaya (myininaya):

ok well lets use the first one as an example 15=3*5 30=3*3*5*5 so the greatest common divisor is 3*5=15 and the product of 15*30 is 15(30) all you have to to find the lcm of two numbers is do their product divided by their gcd here we say 15(30)/15=30 ----------------------------------------- now the other problem we have 15=3*5 20=2*2*5 gcd is 5 there product is 15*20=15(20) so the product divided by the gcd is (15)(20)/5=15(4)=60 so 60 is the least common multiple

OpenStudy (anonymous):

I need two multiples.

myininaya (myininaya):

two multiples of the common multiple?

myininaya (myininaya):

for the second one just do 60times some integer like do 60*2=120 this is a common multiple of 15 and 20

OpenStudy (anonymous):

Also find the greatest common multiple of 36 and 54.

myininaya (myininaya):

wow are you sure?

myininaya (myininaya):

greatest common multiple?

myininaya (myininaya):

thats impossible

OpenStudy (anonymous):

Sorry. find the GCF of35 and 54.

myininaya (myininaya):

ok thats possible lol

myininaya (myininaya):

35=5*7 54=2*3*3*3 write prime factorizations for both like i did here there are no common factors written above for 35 and 54 so we could write 35=1*35 54=1*54 so the gcf of 35 and 54 is 1

OpenStudy (anonymous):

But the answer are 36,30,18 and 72.

myininaya (myininaya):

what are you talking about?

OpenStudy (anonymous):

The answer there are for this problems are 36, 30,18 and 72.

myininaya (myininaya):

the gcf of 35 and 54 is 1

OpenStudy (anonymous):

The answer there are for this problems are 36, 30,18 and 72.

myininaya (myininaya):

then you asked the wrong question the gcf gives you only one number